$SPAD/src/input d02ejf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
xend:SF:= 10.0
n:=3
relabs:="D"
iw:=95
x:SF:=0.0
y:Matrix SF:=
 [[1.0 ,0.0 ,0.0 ]]
tol:=0.0001
ef:Expression Float:=(Y[1]::EXPR FLOAT)-0.9
g:Asp9(G):=retract ef
vef:Vector Expression Float:=
 [-0.04*Y[1]+1.0E4*Y[2]*Y[3] ,_
   0.04*Y[1]-1.0E4*Y[2]*Y[3]-3.0E7*Y[2]*Y[2] ,_
   3.0E7*Y[2]*Y[2] ]
fcn:ASP7(FCN):=retract vef
pederv:Asp31(PEDERV):=retract vef
vm:Vector MachineFloat:=[2,4,6,8]
output:Asp8(OUTPUT):=coerce vm
result:=d02ejf(xend,# vm,n,relabs,iw,x,y,tol,-1,g,fcn,pederv,output)

———-

References^{[1] nothing}

4